OpenStudy (mrhoola):

Suppose that X is a Poisson random variable with lambda = 9 Approximate the probability that X is less than 6 Without continuity correction

OpenStudy (anonymous):

$$ \Large{ P(X=k) = \frac{e^{-\lambda} \cdot \lambda^k }{k!} ~~,~ k=0,1,2,... \\~\\ P(X\leq 6) = \sum_{k=0}^{6}\frac{e^{-9} \cdot 9^k }{k!} } $$

OpenStudy (anonymous):

any question so far?

OpenStudy (mrhoola):

what does it mean "with continuity correction" I read my text book and I could not find anything mentioning it Ive already check three times

OpenStudy (zarkon):

since they say "approximate the probability" they probably want you to use a normal approximation to the poisson